A swimming pool is 50 m long and 20 m wide and its depth varies uniformly from 1 meter to 3.5 meter. Find the volume of the pool in cubic meters.

Answer by checkley71(8403)

50*20(1+3.5)/2=100*4.5/2=100*2.25=225M^3 IS THE ANSWER.